The gravitational force between two identical spheres is given as 6 x 10-5 N. The spheres are separated by a distance of 0.60 m at the center. What is the mass of each sphere?


1
Expert's answer
2022-03-30T14:31:28-0400
"F=G\frac{m^2}{r^2}"

"m=\sqrt{Fr^2/G}=\\
\sqrt{6*10^{-5}*0.60^2/6.67*10^{-11}}=569\:\rm kg"